Mal87:

Completely factor the polynomial below. 8q^2+6q-5 A. (4q-1)(2q+5) B. (4q+1)(2q-5) C. (4q-5)(2q+1) D. (4q+5)(2q-1)

Hero:

@Mal87 let's see your steps for this.

Mal87:

We have, => 8q² + 6q - 5 = 0 By middle term splitting method => 40 = 4 x 10 Thus => 8q² + 10q - 4q - 5 = 0 => 2q(4q + 5) - (4q + 5) = 0 => (D) (4q + 5)(2q - 1)

Mal87:

@Hero

Hero:

That's the best response I've seen all day. Looks like you figured it out quicker than I thought you would. At least I think you did. Can't look over your shoulder. I notice though you like skipping steps. Just to make sure you have it see if you can factor \(5x^2 + 23x + 12\)